OpenStudy (anonymous):

find the formula of a linear and a power function passing through the given points: (1,-7) and (3,-1701)

OpenStudy (triciaal):

linear function same as line given points: (1,-7) and (3,-1701) you can find the slope this is the change in y/corresponding change in x m = (Y2 - y 1)/(x 2 - x 1) then using one of the given points and any point (x, y) find the equation of the line

OpenStudy (anonymous):

what about power function?

OpenStudy (triciaal):

find the formula of a power function passing through the given points: (1,-7) and (3,-1701) I am thinking that it would be a circle with the radius being the line joining the 2 points

OpenStudy (anonymous):

power functions is the only one I'm struggling with right now cause I missed the lecture for it. Would you mind explaining to me further the equation of that if you don't mind?

OpenStudy (triciaal):

it could be a parabola with the vertex at (1,- 7)?

OpenStudy (triciaal):

power function i think exponent 2 or more equation of a circle x^2 + y^2 = R^2 where (x,y) lies on the circle with center the origin(0,0) and the radius - r parabola ax^2 + bx + c = 0